i have this question t(n)=4n-3 and it says write down the first five terms of the sequence. i do not understand how to find the terms of this sequence.
Answer by scott8148(6628)

n represents the number of the term __ so just plug in 1 through 5 for n

t(1)=4(1)-3=4-3=1 __ t(2)=4(2)-3=8-3=5 __ etc.